MS1 Combinatorial Matrix Theory

Organizers: Leslie Hogben, Iowa State University, and Bryan Shader, University of Wyoming.

MS1 Wednesday Morning

• 10:35 A Conjecture in Combinatorial Matrix Theory, Brualdi, Richard A., University of Wisconsin, Madison, USA

• 11:00 On nonnegative sign equivalent and sign similar factorizations of matrices Hershkowitz, Danny, Technion, Israel (with Allan Pinkus)

• 11:25 Constructing Laplacian Integral Split Graphs, Kirkland, Steve, University of Regina, Canada, (with N. Abreu, M. de Freitas and R. Del Vecchio)

• 11:50 Potential Diagonal and D-convergence, Pryporova, Olga, Iowa State University, Ames, IA, USA

• 12:15 The Colin de Verdière Parameter- a progress report, Berman, Avi, Technion, Tel Aviv, Israel

MS1 Thursday Morning

• 10:35 Minimum rank of edge subdivisions of complete graphs and wheels, Barrett, Wayne, Brigham Young University, Utah, (with Ryan Bowcutt, Mark Cutler, Seth Gibelyou, and Kayla Owens)

• 11:00 Computing the minimum rank of partial 2-trees, van der Holst, Hein, Eindhoven University of Technology, The Netherlands

• 11:25 Characterizing graphs with minimum rank at most a given number over a finite field using polarities of projective geometries, Grout, Jason, Iowa State University, Ames, IA, USA

• 11:50 Average minimum rank of a graph, Shader, Bryan, University of Wyoming, USA, (with Francesco Barioli, Shaun Fallat, Tracy Hall, Daniel Hershkowitz, Leslie Hogben, Ryan Martin, and Hein van der Holst)

• 12:15 The graph and rank of a positive semidefinite matrix, Deaett, Louis, University of Wisconsin, Madison, USA

MS1 Friday Afternoon

• 15:30 Group Inverses of Matrices with Path Graphs, Olesky, Dale, University of Victoria, Canada, (with M. Catral and P. van den Driessche)

• 15:55 Sparse Inertially Arbitrary Sign Patterns, Vander Meulen, Kevin, Redeemer College, Canada, (with L. Vanderspek and M. Cavers)

• 16:20 The Q-matrix completion problem , DeAlba, Luz, Drake University, Des Moines, IA, USA, (with Leslie Hogben and Bhaba Sharma)

• 16:45 Maximizing spectral radius of unoriented Laplacian matrix, Tam, Bit-Shun, Tamkang University, Tamsui, Taiwan, (with Ding-Jung Chang and Shui-Hei Wu)

MS2 Eigenproblems: Theory and computation

Organizer: Julio Moro, Universidad Carlos III de Madrid

MS2 Thursday Afternoon

• 16:55 On some modified root-finding problems , Dhillon, Inderjit, University of Texas, Austin, TX, USA, (with Matyas Sustik)

• 17:20 Eigen-reduction of Large Scale Neuronal Networks , Cox, Steve, Rice University, Houston TX, USA, (with Tony Kellems, Derrick Roos and Nan Xiao)

• 17:45 Recycling Ritz vectors in the parameterized Lanczos method, Meerbergen, Karl, Katholieke Universiteit Leuven, Belgium, (with Zhaojun Bai)

• 18:10 Numerical methods for two-parameter eigenvalue problems, Plestenjak, Bor, University of Ljubljana, Slovenia

• 18:35

MS2 Friday Morning

• 10:35 Lipschitz stability of canonical Jordan bases of H-selfadjoint matrices, Olshevsky, Vadim, University of Connecticut, Storrs, CT, USA, (with Tom Bella and Upendra Prasad)

• 11:00 Eigenproblems for quasiseparable matrices, Bella, Tom, University of Connecticut, Storrs, CT, USA, (with Yuli Eidelman, Israel Gohberg and Vadim Olshevsky)

• 11:25 Eigenvalues of the sum of matrices from unitary similarity orbits, Li, Chi-Kwong, College of William and Mary, Williamsburg, VA, USA, (with Yiu-Tung poon and Nung-Sing Sze)

• 11:50 Pseudospectra and Stability radii for Hamiltonian Matrices, Karow, Michael, Technische Universitat Berlin, Germany

• 12:15 Structured Holder condition numbers for eigenvalues under fully nongeneric perturbations, Moro, Julio, Universidad Carlos III de Madrid, Spain, (with María J. Peláez)

MS3 Implementation and Application issues in regularizing least squares and total least squares

Organizer: Rosie Renaut, Arizona State University

MS3 Thursday Afternoon

• 16:55 A Truncated Singular Component Method for Ill-Posed Problems, Bert W. Rust, NIST, USA, (with Dianne P. O’Leary, University of Maryland)

• 17:20 Truncation Rules for Iterative Deblurring Methods, John Bardsley, University of Montana, USA

• 17:45 Lanczos Hybrid Regularization for Image Processing Applications, James Nagy, Emory University, Atlanta GA, USA

• 18:10 On solvability of total least squares problem, Iveta Hnetynková, Charles University in Prague, and Institute of Computer Science, Academy of Sciences, Czech Republic.

• 18:35 A Newton Iteration for estimating the regularizing parameter for least squares, Rosemary A Renaut, Arizona State University, Phoenix, USA, (with Jodi Mead, Boise State University)

MS3 Friday Morning

• 10:35 Block splitting least square regularization for structured matrices arising in nonlinear microwave imaging, Claudio Estatico, University of Cagliari, Italy

• 11:00 Signal and Image regularization via antireflective transform, Antonio Aricó, Università degli studi di Cagliari, taly

• 11:25 Regularizing Least Squares with the Concentration Method, Wolfgang Stefan, (with Rosemary Renaut and Anne Gelb, Arizona State University)

• 11:50 Calculating Weights in Least Squares Estimation Using the Chi-squared Method, Jodi Mead, Boise State University, Idaho, (with Rosemary Renaut, ASU)

• 12:15

MS4 Linear Algebra Education

Organizers: María Trigueros, Instituto Tecnológico Autónomo de México, and Asuman Oktac c, Centro de Investigación y Estudios Avanzados, IPN, México, and PUCV, Chile.

MS4 Monday Morning

• 10:45 Intellectual Need and Its Role in Mathematics Instruction: Focus on Linear Algebra, Harel, Guershon, University of California, San Diego, USA

• 11:10 Thinking Modes Revealed on Students’ Responses from an Assignment on Linear Independence, Dogan-Dunlap, Hamide, University of Texas at El Paso, USA

• 11:35 Basic notions of Vector Space Theory: students’ models and conceptions, Maracci, Mirko, Siena University, Italy

• 12:00 Construction of a vector space schema, Parraguez, Marcela, PUCV, Chile, (with Oktac c, Asuman)

• 12:25 Spanning sets and vector spaces they generate: an APOS analysis, Trigueros, María, Instituto Tecnológico Autónomo de México, (with Ku, Darly and Oktac c, Asuman)

MS4 Tuesday Afternoon

• 16:55 Few pedagogical scenarios in Linear Algebra with Cabri and Maple, Klasa-Bompoint, Jacqueline, Dawson College, Canada

• 17:20 When MATLAB Gives ”Wrong” Answers, Leon, Steven, University of Massachusetts Dartmouth, USA

• 17:45 A fundamental paradox in learning algebra, Tanguay, Denis, UQAM, Montreal, Canada, (with Corriveau, Claudia)

• 18:10 Integrating learning theories and application-based modules in teaching linear algebra, Martin, William, North Dakota State University, (with S. Loch, L. Cooley, S. Dexter and D. Vidakovic)

• 18:35 Group discussion

MS4 Wednesday Morning

• 10:35 Skills, Concepts and Models in a Linear Algebra Course, Moura, Ana, Instituto Superior Técnico, UTL, Lisbon, Portugal, (with Santos, P. A.)

• 11:00 Design of a unit to teach eigenvectors and eigenvalues based on the instructional design principles of Realistic Mathematics Education, Zandieh, Michelle, Arizona State University, (with Rasmussen, C. and Larson, C.)

• 11:25 Use of models in the teaching of linear algebra, Possani, Edgar, Instituto Tecnológico Autónomo de México, (with Preciado, G. and Lozano, D.)

• 11:50 A Java Applet and Introductory Tutorial for the Jacobi, Gauss-Seidel and SOR Methods, Strong, David, Pepperdine University, CA, USA

• 12:15 Group discussion

MS5 Linear Algebra in Model Reduction

Organizer: Peter Benner, Chemnitz University of Technology, Germany

MS5 Thursday Morning

• 10:35 A Krylov-Based Descent Approach for the Optimal H2 Model Reduction of Large-Scale Dynamical Systems, Gugercin, Serkan, Virginia Tech, Blacksburg, VA, USA

• 11:00 H2 Approximation and Tangential Rational Interpolation, Van Dooren, Paul, Université Catholique Louvain-La-Neuve, Belgium • 11:25 Interpolatory Projection Methods for Parameterized Model Re-

duction, Beattie, Christopher, Virginia Tech, Blacksburg, VA, USA

• 11:50 Parametric Model Reduction for Systems with Coupled Parameters, Feng, Lihong, Chemnitz University of Technology, Germany/Fudan University Shanghai, China

• 12:15 The Effects of Deflation in Projection-Based Order Reduction, Freund, Roland, University of California, Davis, USA

MS5 Friday Afternoon

• 15:30 Model Reduction for unstable Systems based on Hierarchical Matrix Arithmetic, Baur, Ulrike, Chemnitz University of Technology, Germany

• 15:55 Algebraic Gramians and Model Reduction for Different System Classes, Damm, Tobias, TU Kaiserslautern, Germany

• 16:20 Balancing-Related Model Reduction for Large-Scale Unstable Systems, Benner, Peter, Chemnitz University of Technology, Germany

MS6 Matrix functions and matrix equations

Organizers: Peter Lancaster, University of Calgary, Canada, and Chun-Hua Guo, University of Regina, Canada

MS6 Tuesday Morning

• 10:35 Fast solution of a certain Riccati Equation through Cauchylike matrices, Bini, Dario, University of Pisa, Italy, (with Beatrice Meini and Federico Poloni)

• 11:00 On the numerical solution of large-scale sparse discrete-time Riccati equations, Fassbender, Heike, TU Braunschweig, Germany, (with Peter Benner)

• 11:25 On Newton’s method and Halley’s method for p-th roots of matrices, Guo, Chun-Hua, University of Regina, Canada

• 11:50 Matrix iterations for matrix functions, Iannazzo, Bruno, Università dell’Insubria, Como, Italy

• 12:15 Tensor Sylvester matrices and information matrices of multiple stationary processes , Klein, Andre, University of Amsterdam, The Netherlands

MS6 Tuesday Afternoon

• 16:55 Linearization of Matrix Polynomials, Lancaster, Peter, University of Calgary, Canada

• 17:20 From algebraic Riccati equations to unilateral quadratic matrix equations: old and new algorithms, Meini, Beatrice, University of Pisa, Italy, (with Bini, Dario and Poloni, Federico)

• 17:45 Exponential Integrators for Solving Large-Scale Differential Riccati Equations, Mena, Hermann, Escuela Politiécnica Nacional, Quito, Ecuador, (with Benner, Peter)

• 18:10 Matrix polynomials, rational matrices and linear systems: A review, Zúñiga, Juan Carlos, University of Calgary, Canada, (with Lancaster, Peter)

• 18:35

MS7 Max Algebra

Organizers: Hans Schneider, University of Wisconsin, and Peter Butkovic, University of Birmingham, UK

MS7 Monday Morning

• 10:45 Using max-plus eigenvalues to bound the roots of a polynomial, Gaubert, Stephane, INRIA, Paris, France, (with Marianne Akian (INRIA)and Adrien Brandejsky (ENS Cachan))

• 11:10 Bounds for the Perron root using max eigenvalues, van den Driessche, Pauline, University of Victoria, Canada, (with Elsner, Ludwig)

• 11:35 Max-Plus Bases, Cornices and Pruning, McEneaney, William, University of California, San Diego, USA

• 12:00 Solving two-sided (max,plus)-linear equation systems, Zimmermann, Karel, Charles University, Prague, Czech Republic

• 12:25 NP-hard problems in extremal algebras tackled by particle swarm optimization, Ponce, Daniela, University of Hradec Kralove, Czech Republic, (with Gavalec, Martin)

MS7 Tuesday Afternoon

• 16:55 On Steady State Controller in Min-Plus Algebra, Boimond, Jean-Louis, University Angers, France, (with S. Lahaye)

• 17:20 On Kleene stars and intersection of finitely generated semimodules, Sergeev, Sergey, University of Birmingham, UK

• 17:45 Semi-group of matrices acting on the max-plus projective space, Merlet, Glenn, LIAFA CNRS-Universite Paris-Diderot, France

• 18:10 On the permuted max-algebraic eigenvector problem, Butkovic, Peter, University of Birmingham, UK

• 18:35

MS7 Wednesday Morning

• 10:35 Max-min convexity, Singer, Ivan, Institute of Mathematics, Bucharest, Romania

• 11:00 Robust stability of positive difference equations, Loiseau, Jean Jacques, University Nantes, France, (with M. Di Loreto)

• 11:25 Permuted max-min eigenvector problem, Gavalec, Martin, University of Hradec Kralove, Czech Republic, (with J. Plavka)

• 11:50 On the robustness of matrices in max-min algebra, Plavka, Jan, Technical University Kosice, Slovakia

• 12:15 Nonnegative linear algebra and max linear algebra: where’s the difference?, Schneider, Hans, University of Wisconsin, Madison, USA

MS8 Nonnegative and eventually nonnegative matrices

Organizer: Judi McDonald, Washington State University, USA

MS8 Monday Afternoon

• 16:55 Nonnegative and Eventually Nonnegative Matrices, McDonald, Judith, Washington State University, USA

• 17:20 An example of constructing a nonnegative matrix with given spectrum, Smigoc, Helena, University College Dublin, (with Laffey, Thomas)

• 17:45 Some constructive techniques in the nonnegative inverse eigenvalue problem, Laffey, Thomas, University College Dublin, (with Smigoc, Helena)

• 18:10 Generalized Soules Matrices, Eubanks, Sherod, Washington State University, USA • 18:35 The Kemeny Constant in Finite Homogeneous Ergodic Markov

Chains, Catral, Minerva, University of Victoria, Canada

MS8 Tuesday Morning

• 10:35 On General Matrices Having the Perron-Frobenius Property, Szyld, Daniel, Temple University, USA, (with Elhashash, Abed)

• 11:00 Jordan forms corresponding to nonnegative and eventually nonnegative matrices, Morris, DeAnne, Washington State University, USA, (with McDonald, Judith)

• 11:25 Convergence of Stationary Iterative Methods for Hermitian Semidefinite Linear Systems, Szyld, Daniel, Temple University, USA, (with Frommer, Andreas and Nabben, Reinhard)

• 11:50 On Optimal Condition Numbers For Markov Chains, Neumann, Michael, University of Connecticut, USA, (with Sze, Nung-Sing)

• 12:15